Line t is parallel to the line y=3x+2 and passes thru the highest point on the graph of y=x(10-x). What is the equation of t?

Thankyou for your help

Answer by scott8148(6628)

parallel means same slope, so y=3x+b

the highest point of the graph is the vertex, which is on the axis of symmetry

y=x(10-x) __ y=-x^2+10x
__ for axis of symmetry, x=-b/2a __ x=-10/(-2) __ x=5

substituting __ y=5(10-5) __ y=25

substituting __ 25=3(5)+b __ 25=15+b __ 10=b

so t is y=3x+10